OpenStudy (anonymous):

Find dy/dx when ln xy- 3x=5

OpenStudy (anonymous):

start with \[\frac{1}{xy}\times (y+xy')-3=0\] and solve for \(y'\)

OpenStudy (anonymous):

or if you prefer, since \(\log(xy)=\log(x)+\log(y)\) you can start with \[\log(x)+\log(y)-3x=5\] then get \[\frac{1}{x}+\frac{y'}{y}-3=0\] and solve for \(y'\) may be quicker

OpenStudy (anonymous):

The answer choices all include e^3x+5 which I dont understand.

OpenStudy (anonymous):

The choices are 1)dy/dx= ((3x-1)/x)e^(3x+5) 2) dy/dx = ((3x-5)/x)e^(3x+5) 3) dy/dx = ((3x-1)/x^2)e^(3x+5) 4) dy/dx = ((1-3x)/x^2)e^(3x+5) 5)dy/dx = ((5-3x)/x)e^(3x+5) 6) dy/dx = ((3x-5)/x^2)e^(3x+5) It is not 1

OpenStudy (lastdaywork):

@saw3266 Can you show your working ??

OpenStudy (anonymous):

what?

OpenStudy (lastdaywork):

Did you use satellite's method ??

OpenStudy (anonymous):

I got y'= -y/x +3y but that doesnt have e^(3x+5) in it

OpenStudy (lastdaywork):

Can you find the value of y from the original quesiton ??

OpenStudy (anonymous):

it is \[\ln(xy)-3x=5\] right? or did i read it wrong?

OpenStudy (anonymous):

i guess you could start with \[\ln(xy)=3x+5\] and so \[xy=e^{3x+5}\] and go from there

OpenStudy (anonymous):

take the derivative of both sides, get \[xy'+y=3e^{3x+5}\] then solve for \(y'\)

OpenStudy (anonymous):

opps. I made a mistake with the answer choices. 2,4,6 are e^3x-5

OpenStudy (anonymous):

OHHH! that make sense! thanks!

OpenStudy (31356):

@satellite73

OpenStudy (anonymous):

i didnt realize the whole ln to e thing. Our teacher is ridiculously smart and skips steps like that.

OpenStudy (31356):

Please help me with this satellite73